OpenStudy (anonymous):

If anyone can help me with a few math questions, I would be very thankful. Please..

OpenStudy (abdullah1995):

go ahead with the questions. i can't promise but i will try my best if i know the answer

OpenStudy (anonymous):

I posted a picture on this post. Do you see it?

OpenStudy (abdullah1995):

no i don't see the picture. try again

OpenStudy (anonymous):

If 3X + Y = 18, then 2(3x + Y) =

OpenStudy (abdullah1995):

what grade are you in ?

OpenStudy (anonymous):

Senior.

OpenStudy (abdullah1995):

well 3x+y = 18 now you multiplied the whole thing by 2(3x+y) in maths, whatever you do to the left side of the equation you have to do the same for the right side of the equation. so 3x+y=18 2(3x+y)=2(18) 2(3x+y)=36

OpenStudy (anonymous):

So the answer is 36 then?

OpenStudy (abdullah1995):

i believe so, yes